# Coupled Input-Output Dimension Reduction: Application to goal-oriented Bayesian experimental design and global sensitivity analysis

### v1

@@ 5.1 Conditioned Diffusion
Figure 5 shows the convergence of the relative $L^{2}$-error and upper bound with increasing subspace dimensions $r$ and $s$. We compare the errors of our coupled dimension reduction method with other dimension reduction methods in the data and parameter spaces such as PCA, CCA [29] and the joint dimension reduction method [12] discussed in Remark 3.12. Our coupled method has very similar error and upper bound values as the joint dimension reduction method. Both decrease much faster with increasing subspace dimensions and do not get stuck on a plateau like CCA.
@@ 5.2 Burgers’ Equation
with $t\in[0,1]$, $x\in[0,1]$, $D=0.001$ and periodic boundaries. Our parameters are the initial conditions $u_{0}(x)$. We assume a normal prior distribution for $u_{0}(x)$ with mean $\mu(x)=\tfrac{1}{\sqrt{2\pi}}\exp(-\tfrac{1}{2}\|0.5-x\|^{2})$ and Matérn covariance with correlation length $l=0.1$ and smoothness $\nu=2.5$. Using a uniform space discretization our parameters are $(u_{0}(x_{1}),\ldots,u_{0}(x_{d}))\in\mathbb{R}^{d}$ with $d=100$. As data $y\in\mathbb{R}^{m}$, we use state observations at the final time $T=0.1$ with i.i.d. measurement noise $\eta\sim\mathcal{N}(0,0.01\,I_{m})$ so that $m=100$. Therefore, $G\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ maps the discretized initial condition $u_{0}(x)$ to the discretized solution $u(x,T)$ at the final time. The left subplot in Figure 6 shows the prior mean $\mu(x)$ together with several prior samples and also visualizes each of the goals $U_{r}^{(i)}$, which project the initial conditions $X$ onto the grid points $i,\ldots,i+r$ marked by an orange bar. The right subplot in Figure 6 illustrates the space-time model solutions $u(x,t)$ with the prior mean as initial condition $u_{0}(x)=\mu(x)$. We consider $r=5$ and the sensor placement for $s=10$ sensors. To compute the diagnostic matrices, we use $M=1000$ prior samples, and the goal-oriented true EIG (23) is computed via the double-loop importance sampling estimator from [14] with $30$ inner and $M$ outer loop samples.
@@
In Figure 7, we present the goal-oriented EIG for different goals $U_{r}^{(i)}$ and sensor placement methods. For each goal, we select sensors based on three different methods: the optimal sensor placement $\tau^{*}$ in (24), the goal-oriented EIM-based sensors $\widetilde{\tau}$ and the PCA-EIM sensors $\widetilde{\tau}^{\textup{PCA}}$. The violin plots also show the EIG values for $100$ random sensor placements. Both sensors $\tau^{*}$ and $\widetilde{\tau}$ almost always outperform the random sensors, i.e., yielding larger EIG. Moreover, the EIM-based sensors $\widetilde{\tau}$ sometimes slightly outperform $\tau^{*}$. On the other hand, the PCA-EIM sensors $\widetilde{\tau}^{\textup{PCA}}$ are often only as good as average random selections, especially for goals $U_{r}^{(i)}$ with lower indices $i<45$. This phenomenon can be explained by looking at the actual sensor positions in Figure 8. Since the solution has a wavefront that is moving to the right, the final state has strong features on the right side of the domain. As a result, the PCA-EIM sensors are concentrated on the right side of the domain, regardless of the goal. Thus, the PCA-EIM sensors seem to be more informative for parameters of interests $U_{r}^{(i)}$ that also lie on the right side of the domain (large $i$) and less so for parameters of interest located on the left side of the domain (small $i$). Figure 8 also shows that the EIM-based indices $\widetilde{\tau}$ are often more spread than the optimal sensors $\tau^{*}$, which could explain their better performance.
Figure 7: Goal-oriented EIG values of different goals $U_{r}^{(i)}$ and sensor placements. Violin plots represent $100$ random sensor placements.
@@ 6 Conclusions
The error analysis associated with such diagnostics is a natural question for future work. Third, our coupled dimension reduction framework can also be extended to operator learning [38, 31] where $G$ is considered as operator $G\colon\mathcal{U}\rightarrow\mathcal{V}$ acting on function spaces $\mathcal{U},\mathcal{V}$, typically separable Hilbert spaces. This extension is possible since the Poincaré and Cramér-Rao inequalities are known to hold in infinite dimensions.
@@ Appendix A Linking the $L^{2}$-error of $\widetilde{G}$
where $\widehat{\pi}_{X|Y}(x|y)\propto\exp(-\tfrac{1}{2\sigma^{2}}\|\widetilde{G}(x)-y\|^{2})\pi_{X}(x)$ satisfies the definition in Proposition A.1 with $f_{1}(x_{r},y_{s})=\tfrac{1}{Z}\exp(-\tfrac{1}{2\sigma^{2}}\|\widetilde{g}(x_{r})+\widetilde{G}_{\perp}-y_{s}\|^{2})$ and $f_{2}(x_{\perp},x_{r})=\pi_{X}(x)$. Here, $Z\in\mathbb{R}$ represents a normalization constant independent of $x$ and $y$. By Proposition 4.1 in [20] we have the following bound for any posterior $\pi_{X|Y}$ with Gaussian likelihood and its approximation $\widehat{\pi}_{X|Y}$ with an approximate forward model
@@ Appendix D Cramér-Rao-like Inequality
$S^{\textup{cl}}(U_{r}|V_{s})$ $=1-\frac{\mathbb{E}\bigl{[}\|V_{s}^{\top}G(X)-\mathbb{E}[\left.V_{s}^{\top}G(X)\right|U_{r}^{\top}X]\|^{2}\bigr{]}}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)}$
$S^{\textup{tot}}(U_{r}|V_{s})$ $=\frac{\mathbb{E}\bigl{[}\|V_{s}^{\top}G(X)-\mathbb{E}[\left.V_{s}^{\top}G(X)\right|U_{\perp}^{\top}X]\|^{2}\bigr{]}}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)}.$
@@
$\mathbb{E}\big{[}\|V_{s}^{\top}G(X)$ $-\mathbb{E}[\left.V_{s}^{\top}G(X)\right|U_{r}^{\top}X]\|^{2}\big{]}$
$\leq\overline{\mathcal{C}}(X)\Big{(}\mathbb{E}\left[\|V_{s}^{\top}\nabla G(X)\|_{F}^{2}\right]-\mathbb{E}\bigl{[}\|V_{s}^{\top}\nabla G(X)U_{r}\|_{F}^{2}\bigr{]}\Big{)}$
$=\overline{\mathcal{C}}(X)\operatorname{Tr}(U_{\perp}^{\top}H_{X}(V_{s})U_{\perp})$
$\mathbb{E}\big{[}\|V_{s}^{\top}G(X)$ $-\mathbb{E}[\left.V_{s}^{\top}G(X)\right|U_{r}^{\top}X]\|^{2}\big{]}$
$\geq\overline{c}(X)\Big{(}\big{\|}V_{s}^{\top}\mathbb{E}\left[\nabla G(X)\right]\big{\|}_{F}^{2}-\left\|V_{s}^{\top}\mathbb{E}\bigl{[}\nabla G(X)\bigr{]}U_{r}\right\|_{F}^{2}\Big{)}$
$=\overline{c}(X)\left\|V_{s}^{\top}\mathbb{E}\left[\nabla G(X)\right]U_{\perp}\right\|_{F}^{2}$
By combining these relations we directly obtain (28) and, by replacing $U_{r}$ with $U_{\perp}$ in the above inequalities, we obtain (29). This concludes the proof.

### v2

@@ 5.1 Conditioned Diffusion
Figure 5 shows the convergence of the relative $L^{2}$-error and upper bound with increasing subspace dimensions $r$ and $s$. We compare the errors of our coupled dimension reduction method with other dimension reduction methods in the data and parameter spaces such as PCA, canonical correlation analysis (CCA) [35] and the joint dimension reduction method [13] discussed in Remark 3.12. Our coupled method has very similar error and upper bound values as the joint dimension reduction method. Both decrease much faster with increasing subspace dimensions and do not get stuck on a plateau like CCA.
@@ 5.2 Burgers’ Equation
with $t\in[0,1]$, $x\in[0,1]$, $D=0.001$ and periodic boundaries. Our parameters are the initial conditions $u_{0}(x)$. We assume a normal prior distribution for $u_{0}(x)$ with mean $\mu(x)=\tfrac{1}{\sqrt{2\pi}}\exp(-\tfrac{1}{2}\|0.5-x\|^{2})$ and Matérn covariance with correlation length $l=0.1$ and smoothness $\nu=2.5$. Using a uniform space discretization our parameters are $(u_{0}(x_{1}),\ldots,u_{0}(x_{d}))\in\mathbb{R}^{d}$ with $d=100$. As data $y\in\mathbb{R}^{m}$, we use state observations at the final time $T=0.1$ with i.i.d. measurement noise $\eta\sim\mathcal{N}(0,0.01\,I_{m})$ so that $m=100$. Therefore, $G\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ maps the discretized initial condition $u_{0}(x)$ to the discretized solution $u(x,T)$ at the final time. The left subplot in Figure 6 shows the prior mean $\mu(x)$ together with several prior samples and also visualizes each of the goals $U_{r}^{(i)}$, which project the initial conditions $X$ onto the grid points $i,\ldots,i+r$ marked by an orange bar. The right subplot in Figure 6 illustrates the space-time model solutions $u(x,t)$ with the prior mean as initial condition $u_{0}(x)=\mu(x)$. We consider $r=5$ and the sensor placement for $s=10$ sensors. To compute the diagnostic matrices, we use $M=1000$ prior samples, and the goal-oriented true EIG (23) is computed via the double-loop importance sampling estimator from [15] with $30$ inner and $M$ outer loop samples making sure the estimator is sufficiently accurate for the subsequent comparison.
@@
In Figure 7, we present the goal-oriented EIG for different goals $U_{r}^{(i)}$ and sensor placement methods. For each goal, we select sensors based on four different methods: the optimal sensor placement $\tau^{*}$ from (24), the goal-oriented EIM-based sensors $\widetilde{\tau}$, the PCA-EIM sensors $\widetilde{\tau}^{\textup{PCA}}$ and sensors $\tau^{F}$ based on the Fedorov-Wynn algorithm with a linearized model around the prior mean composed with the linear goal operator [24, 61]. The violin plots also show the EIG values for $250$ random sensor placements. Both sensors $\tau^{*}$ and $\widetilde{\tau}$ almost always outperform the random sensors, i.e., yielding larger EIG. In general, both $\widetilde{\tau}$ and $\tau^{*}$ perform similarly well as the Fedorov-Wynn-based sensors $\tau^{F}$. Their performance surpasses that of $\tau^{F}$ for the goals $U_{r}^{(45)}$ and $U_{r}^{(55)}$, where non-linear effects may significantly compromise the accuracy of the model linearization in the Fedorov-Wynn algorithm. On the other hand, the PCA-EIM sensors $\widetilde{\tau}^{\textup{PCA}}$ are often only as good as average random selections, especially for goals $U_{r}^{(i)}$ with lower indices $i<45$. This phenomenon can be explained by looking at the actual sensor positions in Figure 8. Since the solution has a wavefront that is moving to the right, the final state has strong features on the right side of the domain. As a result, the PCA-EIM sensors are concentrated on the right side of the domain, regardless of the goal. Thus, the PCA-EIM sensors seem to be more informative for parameters of interests $U_{r}^{(i)}$ that also lie on the right side of the domain (large $i$) and less so for parameters of interest located on the left side of the domain (small $i$). Generally, $\tau^{*}$ and $\widetilde{\tau}$ are similarly distributed as the Fedorov-Wynn-based sensors $\tau^{F}$ except for goal $U_{r}^{(55)}$, where the $\tau^{F}$ exhibits a low EIG. Comparing the EIM-based indices $\widetilde{\tau}$ with the sensors $\tau^{*}$ that optimize the EIG lower bound, we observe in Figure 8 that $\widetilde{\tau}$ tends to be more spread over the domain of high variability in the goal-oriented modes $V_{s}^{*}(U_{r}^{(i)})$. This behaviour could explain their slightly better performance over $\tau^{*}$ relative to the true EIG.
Figure 7: Goal-oriented EIG values of different goals $U_{r}^{(i)}$ and sensor placements. Violin plots represent $250$ random sensor placements.
@@ 6 Conclusions
In particular, a nonlinear output quantity of interest allows for more general goal-oriented BOED problems, while a nonlinear transformation $\psi_{r}$ could map the inputs to independent components (e.g. through a Knothe-Rosenblatt transform) for GSA. The error analysis associated with such diagnostics is a natural question for future work. Third, our coupled dimension reduction framework can also be extended to operator learning [45, 37] where $G$ is considered as operator $G\colon\mathcal{U}\rightarrow\mathcal{V}$ acting on function spaces $\mathcal{U},\mathcal{V}$, typically separable Hilbert spaces. This extension is possible since the Poincaré and Cramér-Rao inequalities are known to hold in infinite dimensions.
Acknowledgements
The authors appreciate discussions of some of the results with Matthew Li. The authors acknowledges support from the ANR JCJC project MODENA (ANR-21-CE46-0006-01). RB also acknowledges support from from the von Kármán instructorship at Caltech, the Air Force Office of Scientific Research MURI on “Machine Learning and Physics-Based Modeling and Simulation” (award FA9550-20-1-0358) and a Department of Defense (DoD) Vannevar Bush Faculty Fellowship (award N00014-22-1-2790) held by Andrew M. Stuart.
@@ Appendix A Linking the $L^{2}$-error of $\widetilde{G}$
where $\widehat{\pi}_{X|Y}(x|y)\propto\exp(-\tfrac{1}{2\sigma^{2}}\|\widetilde{G}(x)-y\|^{2})\pi_{X}(x)$ satisfies the definition in Proposition A.1 with $f_{1}(x_{r},y_{s})=\tfrac{1}{Z}\exp(-\tfrac{1}{2\sigma^{2}}\|\widetilde{g}(x_{r})-y_{s}\|^{2})$ and $f_{2}(x_{\perp},x_{r})=\pi_{X}(x)$. Here, $Z\in\mathbb{R}$ represents a normalization constant independent of $x$ and $y$. By Proposition 4.1 in [23] we have the following bound for any posterior $\pi_{X|Y}$ with Gaussian likelihood and its approximation $\widehat{\pi}_{X|Y}$ with an approximate forward model
@@ Appendix D Cramér-Rao-like Inequality
$S^{\textup{cl}}(U_{\tau}|V_{s})$ $=1-\frac{\mathbb{E}\bigl{[}\|V_{s}^{\top}G(X)-\mathbb{E}[\left.V_{s}^{\top}G(X)\right|U_{\tau}^{\top}X]\|^{2}\bigr{]}}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)}$
$S^{\textup{tot}}(U_{\tau}|V_{s})$ $=\frac{\mathbb{E}\bigl{[}\|V_{s}^{\top}G(X)-\mathbb{E}[\left.V_{s}^{\top}G(X)\right|U_{-\tau}^{\top}X]\|^{2}\bigr{]}}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)}.$
@@
$\mathbb{E}\big{[}\|V_{s}^{\top}G(X)$ $-\mathbb{E}[\left.V_{s}^{\top}G(X)\right|U_{\tau}^{\top}X]\|^{2}\big{]}$
$\leq\overline{\mathcal{C}}(X)\Big{(}\mathbb{E}\left[\|V_{s}^{\top}\nabla G(X)\|_{F}^{2}\right]-\mathbb{E}\bigl{[}\|V_{s}^{\top}\nabla G(X)U_{\tau}\|_{F}^{2}\bigr{]}\Big{)}$
$=\overline{\mathcal{C}}(X)\operatorname{Tr}(U_{-\tau}^{\top}H_{X}(V_{s})U_{-\tau})$
$\mathbb{E}\big{[}\|V_{s}^{\top}G(X)$ $-\mathbb{E}[\left.V_{s}^{\top}G(X)\right|U_{\tau}^{\top}X]\|^{2}\big{]}$
$\geq\overline{c}(X)\Big{(}\big{\|}V_{s}^{\top}\mathbb{E}\left[\nabla G(X)\right]\big{\|}_{F}^{2}-\left\|V_{s}^{\top}\mathbb{E}\bigl{[}\nabla G(X)\bigr{]}U_{\tau}\right\|_{F}^{2}\Big{)}$
$=\overline{c}(X)\left\|V_{s}^{\top}\mathbb{E}\left[\nabla G(X)\right]U_{-\tau}\right\|_{F}^{2}$
By combining these relations we directly obtain (28) and, by replacing $U_{\tau}$ with $U_{-\tau}$ in the above inequalities, we obtain (29). This concludes the proof.